OpenStudy (anonymous):

If x*x + 2x - 35 = 0, then x = ?

OpenStudy (anonymous):

will give medals

OpenStudy (anonymous):

@sexythang101 could u help

OpenStudy (anonymous):

x^2 +7x-5x-35=0 (x+7)(x-5) =0 x=-7 and 5

OpenStudy (anonymous):

thnx @gyanu

OpenStudy (anonymous):

you are welcome @nova2528